Sheila has just arrived at the airport and is dragging her suitcase to the luggage check-in desk. She pulls on the strap with a force of 190 N at an angle of 35° to the horizontal to displace it 45 m to the desk. If the mass of her suitcase is 20 kg ,what is the speed of the suitcase when it reaches the check -in desk.
Answer #1

apply work energy therom

work done by Sheila on suitcase = change of KE of suitcase

F.D*cosθ =1/2* m*( vf^2-vi^2)

190*45*cos35 = 0.5*20*(vf^2 -0)

from here vf = 26.4646 m/s

let me know in a comment if there is any problem or doubts​